OpenStudy (anonymous):

-3x - 8 + 6x = 7 solve for x

OpenStudy (anonymous):

x = -5

OpenStudy (anonymous):

are you sure its not x=5?

OpenStudy (anonymous):

3x=15 x=5

OpenStudy (anonymous):

taking x terms to LHS and constants to RHS \[-3x + 6x = 7+8\] \[3x = 15\] \[x = 5\]

OpenStudy (anonymous):

thanks